A chemist mixed some 15% peroxide solution with some 6% peroxide solution to make 300ml of a 9% peroxide solution. How many milliliters of the 15% solution were used?
Answer by nitin.thacker(3)

Say there is x ml of 15% concentration peroxide and there is y ml of 6% peroxide. This mixture makes 300 ml of 9% concentration peroxide.
So our equations becomes
x + y = 300
and
15*x + 6*y = 9*300
Solving this gives x = 100 and y =200.
So the answer is 100 ml of 15% peroxide is used to make solution.